If a consumer's objective function is given by U=X1X2, subject P1X1+P2X2=M established the demand function for X1 and X2 that maximize the consumer satisfaction
1
Expert's answer
2021-09-23T09:14:02-0400

Consumer get maximum satisfaction when his/her utility maximize

Given

"U=X_1X_2, \\subject \\P_1X_1+P_2X_2=M"

Using langrarian we can maximize utility

"L=X_1X_2+\lambda[M-P_1X_1-P_2X_2]
\\\frac{\delta L}{\delta X_1}=X_2-\lambda P_1=0.....(1)\\
\frac{\delta L}{\delta X_2}=X_1-\lambda P_2=0.....(2)\\

\frac{\delta L}{\delta \lambda}=M-P_1X_1-P_2X_2=0.....(3)"


Divide equition 1 by 2

"\frac{X_2}{X_1}=\frac{\lambda P_1}{\lambda P_2}=\frac{X_2}{X_1}=\frac{P_1}{P_2}\\P_2X_2=P_1X_1" Put in equition 3


"M=P_1X_1+P_2X_2\\M=P_1X_1+P_1X_1"

Therefore the demand of X1 and X2 that will maximize consumer satisfaction will be

"X_1=\frac{M}{2P_1}\\X_2=\frac{M}{2P_2}"